OpenStudy (anonymous):

in a farm there are 600 banana trees. the number of trees in each row is 10 less than twice the number of rows. how many trees are in each row?

OpenStudy (colcaps):

x (2x - 10) = 600 2x² - 10x - 600 = 0 x² - 5x - 300 = 0 x = 20

OpenStudy (colcaps):

20 trees in each row

OpenStudy (mayankdevnani):

Let there are x rows. Then number of trees in each row is given by (2x - 10). Hence the total number of trees is x (2x - 10) = 600 or, 2x² - 10x - 600 = 0 or, x² - 5x - 300 = 0 It solves to x = 20. So there 20 rows and 30 trees in each row.

OpenStudy (mayankdevnani):

@RDB

OpenStudy (anonymous):

20 rows, 30 trees